Fixed costs $280,000

Variable cost per procedure $80

Volume 10,000 visits

Given the information above:

  1. What revenue per visit is required to break even?
  2. What revenue per visit is required to provide a profit of $100,000?
  3. Using the answer to question 2 above, find the contribution margin required to provide a profit of $100,000
Answer #1


Solution: a) Profit=(SP-VC)*Q-FC 0=SP-80)*10000-280000 SP=1280000/10000)+80 SP= 108 b) Profit=(SP-VC)*Q-FC 100000=(SP-80)*100